OpenStudy (anonymous):

Will give Medal!Ronnie had a certain number of video games, v. One-third of those games were role-playing games. He bought twelve more role-playing games to reach a grand total of twenty-nine role-playing games. What was the original number of video games Ronnie had? 34 51 48 17

OpenStudy (anonymous):

before he bought the12 role playing games to get a total of 29, how many role playing games did he have?

OpenStudy (anonymous):

put another way, what is \(29-12\) ?

OpenStudy (anonymous):

17

OpenStudy (anonymous):

ok good, and that was one third of the games he started with

OpenStudy (anonymous):

17 is one third of what number?

OpenStudy (anonymous):

Sorry not good with fractions

OpenStudy (anonymous):

lets put it another way then what is \(3\times 17\) ?

OpenStudy (anonymous):

51

OpenStudy (anonymous):

ok that looks good

OpenStudy (anonymous):

not sure what "not good with fractions" has to do with anything one third of a number is the same as dividing an number by three can't be that hard can it?

OpenStudy (anonymous):

Thank you so much! I am working this on a tablet not fun.

OpenStudy (anonymous):

True it just not easy for me and I thought it was fraction

OpenStudy (anonymous):

fraction is just a number like any other

OpenStudy (anonymous):

Can you answer another question?

OpenStudy (anonymous):

depends on the question

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

Cormac is making 27 phone calls each week as part of a fundraiser. Last week, he made 11 phone calls on weekdays. Which equation could be used to find x, the number of phone calls he made last week on weekend days? x - 27 = 11 x + 27 = 11 27x = 11 11 + x = 27

OpenStudy (anonymous):

@satellite73